1. Consider the market supply curve which passes through the intercept and from which the market equilibrium data is known, this is, the price and quantity of equilibrium 𝑷𝑬=πŸ“πŸŽ and 𝑸𝑬=𝟐𝟎𝟎𝟎.

a. Considering those two points, find the equation of the supply.

b. Draw a graph of this line.


2. Considering the previous supply line, determine if the following demand function corresponds to the market demand equilibrium stated above. 𝑸𝑫=πŸ‘πŸŽπŸŽπŸŽβˆ’πŸπ’‘.


3. The production function of a firm is described by the following equation 𝑸=𝟏𝟎,πŸŽπŸŽπŸŽπ‘³βˆ’πŸ‘π‘³ 𝟐 where L stands for the units of labor.

a) Draw a graph for this equation. Use the quantity produced in the y-axis and the units of labor in the x-axis.

b) What is the maximum production level?

c) How many units of labor are needed at that point?


4. Solve the following system of equations.

50π‘₯+20𝑦=1800 10π‘₯+3𝑦=300




1
Expert's answer
2022-03-22T19:05:15-0400

The equation is (2000,50)

To find the slope we do as follows

"m = (0,0),(2000,50)"

"m= \frac{1}{40}"

"\frac{1}{40} = (2000,50), (Q,P)"

"40P -2000= Q-2000"

"P= \frac{Q}{40}" or Q=40P

The above is the supply equation.

The graph of this line is as shown below



Considering the previous supply line;

"Q=4P"


To find equilibrium, we equate it to the given Qd.

"40P = 3000-2P"

"P= 71.42"

From this result the Qd does not correspond to the market equilibrium.


"50x+20y= 1800"

"10x+3y=300"

Solving the equations simultaneously

Multiply equation 2 by 5

"50x +20y = 1800"

"50x+ 15y = 1500"

"5y =300"

"y = 60"

Solving for x

"50x+1200 =1800"

"50x= 600"

"x = 12"